OpenStudy (anonymous):

im lookin for x an y intercepts for -2x+5y=-15

OpenStudy (karatechopper):

for y int make x=0 and solve for x int make y=0 and solve

OpenStudy (anonymous):

First start with the definitions. What is x-intercept?

OpenStudy (anonymous):

im lookin it over

OpenStudy (anonymous):

Good. Google is good for that kind of stuff.

OpenStudy (karatechopper):

lol

OpenStudy (anonymous):

lost

OpenStudy (karatechopper):

ok well.. x intercept=The x-intercept of a line is the point at which the line crosses the x axis. ( i.e. where the y value equals 0 ) x-intercept = ( x, 0 ) y intercept=The y-intercept of a line is the point at which the line crosses the y axis. ( i.e. where the x value equals 0 ) y-intercept = ( 0, y )

OpenStudy (anonymous):

x intercept: -2x+5(0)=-15 -2x=-15 x=15/2---->(15/2,0) y intercpet: -2(0)+5y=-15 5y=-15 y=-3---->(0,-3)

OpenStudy (anonymous):